A sample of nitrogen gas in a 1.77-L container exerts a pressure of 1.34 atm at 35 degrees C. what is the pressure if the volume of the container is maintained constant and the temp is raised to 343 degrees C.
Answer #1

Given:
Pi = 1.34 atm
Ti = 35.0 oC
= (35.0+273) K
= 308 K
Tf = 343.0 oC
= (343.0+273) K
= 616 K

use:
Pi/Ti = Pf/Tf
1.34 atm / 308.0 K = Pf / 616.0 K
Pf = 2.68 atm
Answer: 2.68 atm